OpenStudy (anonymous):

Two groups of students were asked how far they lived from their school. The table below shows the distances in miles: Group A (distance in miles) 1 1.5 8 9.2 6.8 4.5 4.8 2.5 0.76 Group B (distance in miles) 2 2.5 3.23 1.3 1.8 2.4 3 1.5 1.8 Which statement best compares the mean distances for the two groups? A. It is equal for Group A and Group B. B. It is greater for Group B than Group A. C. Its value for group A is double the value for Group B. D. Its value for group B is three times the value for Group A.

OpenStudy (secret-ninja):

C.? Just a guess. :]

OpenStudy (envy_seven):

For those looking at this and who want an answer I believe it's B. Group A) 1, 1.5, 1.5, 1.8, 2.2, 2.5, 2.8, 3, 3.2 = 19.5/9 = 2.2 Group B) 1.5, 1.8, 1.8, 2, 2.4, 2.5, 3, 3, 3.2 = 21.2/9 = 2.4 The mean for Group B is .2 higher than Group A